OpenStudy (anonymous):

Solve for x: -3 (x + 7) = -12 A. x= 5/3, x= 19/3. B. x= -3, x= -11. C. x= -3, x= 11. D. No solution.

OpenStudy (anonymous):

1st, divide each side by -3 Then plug some of your answer choices in to see if they work

OpenStudy (anonymous):

@mthompson440 I still don't get it.

OpenStudy (anonymous):

What do you get when you divide -12 by -3?

OpenStudy (anonymous):

@mthompson440 4.

OpenStudy (anonymous):

yes, so your new equation now looks like x+7=4 So what plus seven equals four?

OpenStudy (anonymous):

@mthompson440 -3

OpenStudy (anonymous):

yes, -3 is one of your answers. Good job. Now, what else? What number could you take seven away from and still have four?

OpenStudy (anonymous):

I'm sorry I meant what number could you take seven away from and have -3?

OpenStudy (anonymous):

@mthompson440 11?

OpenStudy (anonymous):

Yes, but think positive seven = -3 -11+7 = -3

OpenStudy (anonymous):

wow, my typing is bad

OpenStudy (anonymous):

Would it be x= -3 and x= 11?

OpenStudy (anonymous):

-11+7=4 so you answers are -3, and -11

OpenStudy (anonymous):

Thank you so much. @mthompson440